Circular No. 8686

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2005 P 1 AND S/2005 P 2 S. A. Stern, Southwest Research Institute (SwRI); H. A. Weaver, Ap-

plied Physics Laboratory, Johns Hopkins University; M. J. Mutchler, Space Telescope Science Institute; A. J. Steffl, W. J. Merline, and J. R. Spencer, SwRI; M. W. Buie, Lowell Observatory; and E. F. Young and L. A. Young, SwRI, report color observations of the newly discovered small satellites of Pluto (cf. IAUC 8625, 8676) using the Hubble Space Telescope (+ ACS/HRC) on Mar. 2.75 UT. During these observations, S/2005 P 1 was at an average position of 2′′.69 from the photo-center of Pluto at p.a. 138^{o}, and S/2005 P 2 was 1′′.43 away at p.a. 214^{o}. Using ACS filters F606W and F435W, and adopting for Pluto I (Charon) the value B − V = +0.710 (Buie et al. 1997, Icarus 125, 233), B–V colors +0.653 ± 0.026 and +0.654 ± 0.065 were found for S/2005 P 1 and S/2005 P 2, respectively. These colors indicate that the observed hemispheres of all three of Pluto’s satellites are essentially neutral, with only minor differences among them, as might be expected for bodies having a similar origin and sharing a similar environment. Given that a much redder color of B − V = +0.91 ± 0.15 was reported for S/2005 P 2 (Buie et al. 2006,

http://xxx.lanl.gov/abs/astro-ph/?0512491), additional observations

are solicited.

SUPERNOVAE 2006aq, 2006ar, AND 2006as CBET 423 contains details of a type-Ia supernova (mag ≈ 19.9, cali-

brated to R) discovered by the “Nearby Supernova Factory” collaboration on NEAT images taken on Feb. 24.5 UT: SN 2006as is located at α = 13^{h}51^{m}24^{s}.88, δ = −11^{o}15′26′′.8 (equinox 2000.0). They also obtained spectra (range 320–1000 nm) of supernovae 2006aq and 2006ar (cf. IAUC 8684) on Mar. 7.6 and 7.5 UT, respectively; 2006aq is a type-II event that appears to have faded rapidly relative to the reported discovery magnitude. Type-Ia supernova 2006ar is before or near maximum.

COMET C/2006 A1 (POJMANSKI) Visual total-magnitude estimates: Feb. 11.34 UT, 6.2 (L. A. Mansilla,

Rosario, Argentina, 7×50 binoculars); 27.32, 5.2 (W. Souza, Mairipora, Brazil, naked eye); Mar. 3.83, 5.1 (K. Yoshimoto, Yamaguchi, Japan, 7×35 binoc.); 8.21, 5.4 (J. J. Gonzalez, Leon, Spain, 7×50 binoc.; 2^{o}.5 tail in p.a. 285^{o}); 9.20, 5.8 (A. Baransky, Pylypovychi, Ukraine, 10×50 binoculars).

2006 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green